a 40mm hole is drilled at a speed of 30 m/min and feed of 0.1 mm tooth calculate feed per minute


1
Expert's answer
2021-06-24T21:49:01-0400

"D= 0.04 m"


"V = 30 \frac {m}{min}"


"f = 0.1\times 10^{-4} m"


Cutting Speed "= \frac {\pi \times D \times S} {1000}"


"S =\frac{30 \times 1000}{\pi \times 0.04}=238.73 \times 10^3 \frac {1}{min}"


Feed "= S \times f = 0.1\times 10^{-4} \times 238.73 \times 10^3 = 2.387" feed per min